# Last Iterate Convergence of Incremental Methods and Applications in Continual Learning

### v1

@@ 3.3 Inexact Proximal Point Evaluations
Combining (3.9) and (3.10) and noticing that ${\bm{x}}_{k-1,T+1}={\bm{x}}_{k}$ and ${\bm{x}}_{k-1,1}={\bm{x}}_{k-1}$, we sum the inequalities over $t\in[T]$ and obtain
$+\frac{1}{2\eta_{k}}\sum_{t=1}^{T}\varepsilon_{k-1,t}^{2}+GT\sum_{t=1}^{T}\varepsilon_{k-1,t}.$
Compared to (3.2) in Lemma 3.3, we now have additional error terms in (3.8) from the inexact evaluations. However, as long as the approximation error at each iteration is small, i.e., solving the subproblem of proximal point evaluation to a certain accuracy, we can still maintain the convergence rate of incremental proximal methods with exact proximal point evaluations. We formalize this in the following theorem.
Theorem 3.6 (Inexact Evaluation).
Under Assumptions 1 and 2, the output ${\bm{x}}_{K}$ of Algorithm 2 satisfies
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\frac{1}{2T\sum_{k=1}^{K}\eta_{k}}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}+\frac{G^{2}T}{2}\sum_{k=1}^{K}\frac{\eta_{k}^{2}}{\sum_{j=k}^{K}\eta_{j}}$
$+\frac{1}{2T}\sum_{k=1}^{K}\sum_{t=1}^{T}\frac{\varepsilon_{k-1,t}^{2}}{\sum_{j=k}^{K}\eta_{j}}+G\sum_{k=1}^{K}\sum_{t=1}^{T}\frac{\varepsilon_{k-1,t}\eta_{k}}{\sum_{j=k}^{K}\eta_{j}}.$
Further, given $\epsilon>0$, if $\sum_{t=1}^{T}\varepsilon_{k-1,t}=\tilde{\mathcal{O}}\big{(}\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{\sqrt{K}}\big{)}$, then there exists a constant step size $\eta$ such that $f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\epsilon$ after $\widetilde{\mathcal{O}}\big{(}\frac{G^{2}T\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{2}}\big{)}$ individual gradient evaluations.
@@
Using Lemma 3.5 and choosing ${\bm{z}}={\bm{z}}_{k-1}$ defined by (2.5) and following the proof of Theorem 3.4 with $w_{k}=\frac{\eta_{K}}{\sum_{j=k+1}^{K}\eta_{j}}$, we arrive at
@@
$+\frac{1}{2T\eta_{K}}\sum_{k=1}^{K}\sum_{t=1}^{T}\varepsilon_{k-1,t}^{2}w_{k-1}+\frac{G}{\eta_{K}}\sum_{k=1}^{K}\sum_{t=1}^{T}\varepsilon_{k-1,t}\eta_{k}w_{k-1}$
@@
$+\frac{1}{2T}\sum_{k=1}^{K}\sum_{t=1}^{T}\frac{\varepsilon_{k-1,t}^{2}}{\sum_{j=k}^{K}\eta_{j}}+G\sum_{k=1}^{K}\sum_{t=1}^{T}\frac{\varepsilon_{k-1,t}\eta_{k}}{\sum_{j=k}^{K}\eta_{j}}.$
To analyze the oracle complexity, we choose the constant step size $\eta_{k}\equiv\eta=\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{GT\sqrt{K}}$ for simplicity and without loss generality, which leads to
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\frac{G\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{2\sqrt{K}}\Big{(}1+\sum_{k=1}^{K}\frac{1}{K-k+1}\Big{)}$
$+\frac{G\sqrt{K}}{2\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}\sum_{k=1}^{K}\frac{\sum_{t=1}^{T}\varepsilon_{k-1,t}^{2}}{K-k+1}+G\sum_{k=1}^{K}\frac{\sum_{t=1}^{T}\varepsilon_{k-1,t}}{K-k+1}.$
If $\sum_{t=1}^{T}\varepsilon_{k-1,t}=\tilde{\mathcal{O}}\big{(}\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{\sqrt{K}}\big{)}$ and thus $\sum_{t=1}^{T}\varepsilon_{k-1,t}^{2}=\tilde{\mathcal{O}}\big{(}\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{K}\big{)}$ (for example letting $\varepsilon_{k-1,t}=\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{T\sqrt{K}}$), we can then guarantee $f({\bm{x}}_{K})-f({\bm{x}}_{*})=\tilde{\mathcal{O}}\Big{(}\frac{G\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{\sqrt{K}}\Big{)}$. Hence, given $\epsilon>0$, to guarantee $f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\epsilon$, the total number of individual gradient evaluations will be
$KT=\tilde{\mathcal{O}}\Big{(}\frac{G^{2}T\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{2}}\Big{)}.$
4 Conclusion
This work provides the first oracle complexity guarantees for the last iterate of standard incremental gradient methods, motivated by applications in continual learning. The obtained complexity bounds nearly match the best known oracle complexity bounds that in the same settings were previously known only for the (uniformly) average iterate. We further provided generalizations of these results to weighted average iterate with increasing weights and variants of methods with permuted ordering of updates. Some questions that merit further investigation remain. In particular, our results for the proximal incremental method – which are most closely related to continual learning settings – require a large value of the regularization parameter to bound the forgetting. Further increasing regularization deteriorates both the bounds on the forgetting and the loss on the current task, so this value in our results can be seen as a “critical value.” However, how the forgetting is affected if the regularization is reduced (which improves the loss on the current task) is not known. This question is also closely related to whether larger theoretically-justified step sizes are possible for incremental gradient/shuffled SGD methods, which has remained open despite renewed interest and advances in theory of shuffled SGD convergence.
Acknowledgments
Supplementary Material
Appendix A Omitted Proofs
See 2.6
Proof.
=\;= $\frac{T(T+1)}{6}\sigma_{*}^{2},$

### v2

@@ B.3 Inexact Proximal Point Evaluations
Using triangle inequalities and decomposing ${\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1}=\eta_{k}\sum_{s=1}^{t}{\bm{g}}_{k-1,s}-\nabla M_{\eta_{k}f_{s}}({\bm{x}}_{k-1,s})+\nabla M_{\eta_{k}f_{s}}({\bm{x}}_{k-1,s})$, we bound the term ${\mathcal{T}}:=\sum_{t=1}^{T}\varepsilon_{k-1,t}\|{\bm{x}}_{k-1,t+1}-{\bm{z}}\|$ in Eq. (B.10) as follows
${\mathcal{T}}\leq\;$ $\sum_{t=1}^{T}\varepsilon_{k-1,t}\big{(}\|{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1}\|+\|{\bm{x}}_{k-1}-{\bm{z}}\|\big{)}$
[…]
$\overset{(i)}{\leq}\;$ $\Big{(}\sum_{t=1}^{T}\varepsilon_{k-1,t}\Big{)}^{2}+2\eta_{k}GT\sum_{t=1}^{T}\varepsilon_{k-1,t}+\frac{1}{4\eta_{k}GT}\sum_{t=1}^{T}\varepsilon_{k-1,t}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2},$
where we use Young’s inequality for $(i)$. Combining Eq. (B.9) and (B.10) with the above bound on ${\mathcal{T}}$ and noticing that ${\bm{x}}_{k-1,T+1}={\bm{x}}_{k}$ and ${\bm{x}}_{k-1,1}={\bm{x}}_{k-1}$, we sum the inequalities over $t\in[T]$ and obtain
$T(f({\bm{x}}_{k})-f({\bm{z}}))\leq\;$ $\frac{1}{2\eta_{k}}\Big{(}1+\frac{1}{2\eta_{k}GT}\sum_{t=1}^{T}\varepsilon_{k-1,t}\Big{)}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\frac{1}{2\eta_{k}}\|{\bm{x}}_{k}-{\bm{z}}\|^{2}$
$+\frac{T(T-1)G^{2}\eta_{k}}{2}+\frac{1}{\eta_{k}}\Big{(}\sum_{t=1}^{T}\varepsilon_{k-1,t}\Big{)}^{2}+3GT\sum_{t=1}^{T}\varepsilon_{k-1,t},$
thus finishing the proof. ∎
See 3.5
@@
Using Lemma B.4 with ${\bm{z}}={\bm{z}}_{k-1}$ defined by Eq. (2.1) and multiplying $\eta_{k}w_{k-1}$ on both sides, we have
$T\eta_{k}w_{k-1}(f({\bm{x}}_{k})-f({\bm{z}}_{k-1}))$
[…]
$+\frac{T(T-1)G^{2}\eta_{k}^{2}w_{k-1}}{2}+\frac{w_{k-1}}{2}\Big{(}\sum_{t=1}^{T}\varepsilon_{k-1,t}\Big{)}^{2}+3GT\eta_{k}w_{k-1}\sum_{t=1}^{T}\varepsilon_{k-1,t}.$
Then we sum the inequalities over $k\in[K]$ and follow the proof of Theorem 3.3. To telescope the terms $\|{\bm{x}}_{k}-{\bm{z}}_{k-1}\|^{2}$, we need $\lambda_{k-1}\leq\frac{1}{1+\frac{1}{2\eta_{k}GT}\sum_{t=1}^{T}\varepsilon_{k-1,t}}$ for $11\leq k\leq K-$ 1 such that
$w_{k-1}\lambda_{k-1}^{2}(1+\frac{1}{2\eta_{k}GT}\sum_{t=1}^{T}\varepsilon_{k-1,t})\leq w_{k-1}\lambda_{k-1}\leq w_{k-2},$
while we maintain other requirements on $\{\lambda_{k}\}$ and $\{w_{k}\}$ to obtain the last iterate convergence as in Theorem 3.3. In particular, we take the same choice that $w_{k}=\frac{\eta_{K}}{\sum_{j=k+1}^{K}\eta_{j}}$ and $\lambda_{k}=\frac{\sum_{j=k+1}^{K}\eta_{j}}{\sum_{j=k}^{K}\eta_{j}}$ for $10\leq k\leq K-$ 1, so it suffices to let $\sum_{t=1}^{T}\varepsilon_{k-1,t}\leq\frac{2\eta_{k}\eta_{k-1}GT}{\sum_{j=k}^{K}\eta_{j}}$. So we arrive at
@@
$+\frac{3G}{\eta_{K}}\sum_{k=1}^{K}w_{k-1}\eta_{k}\sum_{t=1}^{T}\varepsilon_{k-1,t}+\frac{1}{2T\eta_{K}}\sum_{k=1}^{K}\Big{(}\sum_{t=1}^{T}\varepsilon_{k-1,t}\Big{)}^{2}w_{k-1}$
@@
$+3G\sum_{k=1}^{K}\sum_{t=1}^{T}\frac{\varepsilon_{k-1,t}\eta_{k}}{\sum_{j=k}^{K}\eta_{j}}+\frac{1}{2T}\sum_{k=1}^{K}\frac{(\sum_{t=1}^{T}\varepsilon_{k-1,t})^{2}}{\sum_{j=k}^{K}\eta_{j}}.$
Hence, given $\varepsilon>0$ and taking the constant step size $\eta_{k}\equiv\eta=\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{GT\sqrt{K}}$ for simplicity, to maintain the convergence rate as in Theorem 3.3 with inexact proximal point evaluations, it suffices to let $\sum_{t=1}^{T}\varepsilon_{k-1,t}\leq\frac{2\eta GT}{K-k+1}$. Indeed, we have
$3G\sum_{k=1}^{K}\sum_{t=1}^{T}\frac{\varepsilon_{k-1,t}\eta_{k}}{\sum_{j=k}^{K}\eta_{j}}=3G\sum_{k=1}^{K}\frac{\sum_{t=1}^{T}\varepsilon_{k-1,t}}{K-k+1}\leq\frac{5G\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{\sqrt{K}},$
and
$\frac{1}{2T}\sum_{k=1}^{K}\frac{(\sum_{t=1}^{T}\varepsilon_{k-1,t})^{2}}{\sum_{j=k}^{K}\eta_{j}}\leq\frac{2G\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{\sqrt{K}}\sum_{k=1}^{K}\frac{1}{(K-k+1)^{3}}\leq\frac{2.5G\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{\sqrt{K}}.$
It remains to follow the proof of Theorem 3.3, thus finishing the proof. ∎